\begin{thebibliography}{99}

\bibitem{Acharya} R. Acharya: Realization of the Riemann Hypothesis via
Coupling Constant Spectrum, arXIV: math-ph/0803.1818 v5

\bibitem{Borwein}
The Riemann Hypothesis: \textbf{A Resource for the Afficinado and
Virtuoso Alike}. Borwein, Choi, Rooney and Weirathmueller
(Editors), Springer (2008); \\
The theory of the Riemann Zeta-Function, Second Edition, E.C.
Titchmarsh, revised by D. R. Heath-Brown, Clarendon Press, Oxford
(1986).

\bibitem{Khuri}
N. N. Khuri: Inverse Scattering, the Coupling Constant Spectrum
and the Riemann Hypothesis, arXiv: hep-th/0111067 v 1, 7 November
2001; \\
K. Chadan and P. C. Sabatier: Inverse Problems in Quantum Scattering
Theory, Second Edition, Springer-Verlag (1989); \\
R. G. Newton: Inverse Schr\"{o}dinger Scattering in 3 dimensions,
Springer-Verlag (1989).

\bibitem{Levin}
B. J. Levin: Distribution of Zeros of Entire Functions,
American Math. Soc. Traslations of Math monographs, Vol. 5,
Providence, RI, 1964. \\
See also, B. Ghusayni, Entire Functions of Order One and Infinite
Type (preprint).

\bibitem{Footnote} We hasten to correct misstatements in Reference
1. The sentence immediately following Eq.~(30) must be deleted.
Eq.~(40) and Eq.~(41) must be replaced by
\begin{equation}
\alpha = \ln 2, \tag{40}
\end{equation}
\begin{equation}
\beta = 0 \tag{41}
\end{equation}
One arrives at this by setting $\mathcal{X}(0) = \mathcal{X}(1) =
1$ and $\xi(0) = \xi(1) = \frac{1}{2}$.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsmath, amssymb}

\begin{document}

\title{
Further Comments on Realization of Riemann Hypothesis via Coupling Constant Spectrum}

\author{R. Acharya \\
Department of Physics\\
Arizona State University\\
Tempe, AZ 85287-1504}

\maketitle

\begin{abstract}
We invoke \textbf{Carlson's theorem} to justify \textbf{and} to
confirm the results previously obtained on the validity of Riemann
Hypothesis via the coupling constant spectrum of the zero energy
$S$-wave Jost function a la N.~N.~Khuri, for the real, repulsive
inverse-square potential in non-relativistic quantum mechanics in
3 dimensions.
\end{abstract}

In a previous note (hereinafter referred to as I) \cite{Acharya},
we have exhibited the solution for the $S$-wave Jost function at
zero energy (i.e., $k^{2}=0$) for the real, repulsive
inverse-square potential, $V(r)=\frac{\lambda}{r^{2}}$ ($\lambda >
0$). We then argued that it is justified to identify the $S$-wave
Jost function, $\mathcal{X}(s)\equiv F_{+}((s(s-1)$; $k^{2}=0)$ with
the Riemann $\xi(s)$ function \cite{Borwein} (up to an
entire function of order one, with \textbf{No} zeros) where the
coupling constant $\lambda$ is \textbf{constrained} to satisfy the
relation, i.e.,
\begin{equation}\label{eq:1}
\lambda = s(s-1)
\end{equation}

The argument hinged on demonstrating that the $S$-wave Jost
function has \textbf{only} zeros on the line $s_{n} = \frac{1}{2}
+ i \gamma^{'}_{n}$, $n = 1$,~2,~3, \dots $\infty$. This was
accomplished by first demonstrating that the analogue of Eq.~(1.2)
of Khuri's paper \cite{Khuri} (\textbf{without} the presence of the potential
$V(r)$ under the sign of integration), i.e.,
\begin{equation}\label{eq:2}
\big[
    Im\lambda_{n}(i\tau)
\big]
    \int^{\infty}_{0}
\big|
    f(\lambda_{n}(i\tau); \text{ } i\tau; \text{ } r
\big| ^{2} dr = 0
\end{equation}
holds, where the all-important constraint, i.e.,
\begin{equation}\label{eq:3}
    \int^{\infty}_{0}
\big|
    f(\lambda_{n}(i\tau); \text{ } i\tau; \text{ } r
\big|
    ^{2} dr < \infty
\end{equation}
is satisfied, i.e.,
\begin{equation}\label{eq:4}
\big[
    Im\lambda_{n}(i \tau)
\big]
    \frac{1}{\tau} \frac{2}{\pi} \int^{\infty}_{0} r K^{2}_{\nu(i
    \tau)} (r) dr = 0
\end{equation}
where
\begin{equation}\label{eq:5}
\int^{\infty}_{0} r K^{2}_{\nu (i \tau)} (r) dr = \frac{1}{8}
\frac{\pi \nu (i \tau)}{\sin \pi \nu (i \tau)}
\end{equation}
with $\nu (i \tau) = \sqrt{\lambda (i \tau) + \frac{1}{4}}$,
$\nu (i \tau) \neq 1$,~2,~3, \dots $\infty$.
Then, Eq.~(\ref{eq:3}) follows for $\tau \neq 0$ ($\tau > 0$).

Thus,
\begin{equation}
I m \lambda_{n} (i\tau) = 0
\end{equation}
We then followed Khuri and arrived at the conclusion,
\begin{equation}
\lambda_{n} (0) \equiv s_{n} (s_{n} -1)
\end{equation}
for \textbf{all} {\mathversion{bold}$n$}, is \textbf{real} and
\textbf{negative}. When $\lambda_{n}(0)$, for all $n$ is real and
negative, (and by rescaling such that $\lambda_{n} <
-\frac{1}{4}$, $\nu(0) = \sqrt{\lambda_{n}(0)+\frac{1}{4}}$
becomes imaginary and hence R.H.S. of Eq.~(\ref{eq:5}) will become
the ratio of $\frac{\pi \nu(0)}{\sinh \pi \nu(0)}$, at $\tau = 0$.
In other words, the zero energy coupling spectrum lies on the
negative real line for $V(r) = \lambda V^{*}(r)$, $V(r) =
\frac{1}{r^{2}}$, $V^{*} \geqslant 0$. In view of
Eq.~(\ref{eq:1}), i.e., $\lambda = s(s-1)$, one then concluded
that $\mathcal{X}(s)$ has all its infinite number of nonzero zeros
on the ``critical" line, Re $s_{n} = \frac {1}{2}$.

The \textbf{key question} arises whether or not, the positions
$\gamma_{n}$ of the infinite number of non-zero zeros on the
critical line, \textbf{coincide} with the Riemann Hypothesis,
i.e.,

$$
s_{n} = \frac{1}{2} \pm i \gamma_{n}, \text{ } n=1 \text{,~2,~3
\dots} \infty
$$
(Recall that Hardy has proved that $\xi(s)$ has \textbf{infinite number}
of zeros on the critical line.)

In other words, one \textbf{must} demonstrate that the positions
of the zeros obtained from the potential, $V(r) =
\frac{\lambda}{r^{2}}$, i.e.
$$
s^{'}_{n} = \frac{1}{2} \pm i \gamma^{'}_{n}, \text{ } n=1\text{,~2,~3
\dots} \infty
$$
coincide exactly with the positions ($\gamma_{n}$) of the Riemann
zeros. This was \textbf{not} done in our earlier note! \textbf{We propose
to demonstrate this here.} Our tool is Carlson's theorem
\cite{Levin}.

Recall,
\begin{equation}
\mathcal{X}(s^{'}) = e^{\pi s^{'}}.
\end{equation}
and
\begin{equation}
\xi(s^{'}) = C e^{A(s^{'}+1)}
\end{equation}
for real $s^{'} \geqslant 0$, $s^{'} = s-1$, where we are stating
the \textbf{established} fact that \textbf{both} $\mathcal{X}(s)$
and $\xi(s)$ are \textbf{entire} with \textbf{No zeros} for real
$s^{'} > 0$.

Then, we invoke Hadamard's factorization theorem to set
\begin{equation}\label{eq:10}
e^{a+\frac{bs^{'}}{m}} \mathcal{X} (\frac{s^{'}}{m}) - \xi (\frac{s^{'}}{m}) =
0, \quad \frac{s^{'}}{m} = 1, 2, 3 \dots
\end{equation}
Eq.~({\ref{eq:10}) is satisfied iff
\begin{equation}\label{eq:11}
C e^{A} = e^{a} \text{ and } A = b + \pi
\end{equation}

We now \textbf{invoke Carlson's theorem} \cite{Levin}:\\
In order that each entire function $f(z)$ satisfying the
conditions
\begin{align}
f(z) & = 0(1) e^{\alpha|z|} \text{, for some } \alpha < \infty \\
f(iy) & = 0(1) e^{\beta|y|} \text{, for some } \beta < \pi \\
f(n) & = 0 \text{ for each positive integer `$n$'} \\
\text{then } f(z) & \equiv 0 \text{ !}
\end{align}
Thus, we conclude that
\begin{equation}\label{eq:16}
e^{a+\frac{bs^{'}}{m}} \mathcal{X} \bigg( \frac{s^{'}}{m} \bigg) = \xi \bigg( \frac{s^{'}}{m}
\bigg) \text{, \textbf{all} } s^{'}
\end{equation}

Clearly Eq.~(\ref{eq:10}) is satisfied for \textbf{all} $\frac{s^{'}}{m} >
0$, when Eq.~(\ref{eq:11}) holds. An analytic (entire) function
which vanishes in a subdomain of its region of analyticity, \textbf{must}
necessarily vanish in its ``entire" domain of analyticity, i.e.,
for \textbf{all} $s^{'}$! Thus, in a sense, one need not invoke
Carlson's theorem to arrive at Eq.~(\ref{eq:16})!

It is now obvious that the zeros on the critical line of the zero
energy Jost function $\mathcal{X}(s)$ and the zeros (infinite
number, established by G.~H.~Hardy) of Riemann's $\xi(s)$ function
must necessarily coincide, i.e.,
$$
s^{'}_{n} = \frac{1}{2} \pm i \gamma^{'}_{n} = s_{n} = \frac{1}{2}
\pm i \gamma_{n}, \quad n=1,2, \dots \infty.
$$

We can explicitly verify this assertion as follows:

We invoke the Hadamard
factorization for every entire function $g(z)$ of order one and
``infinite type" (which guarantees the existance of \textbf{infinitely
many} non-zero zeros [valid for \textbf{all} $z$!] \cite{Levin}:
\begin{equation}\label{eq:17}
g(z) = z^{m} e^{B} e^{Dz} \prod^{\infty}_{n=1} \bigg( 1 -
\frac{z}{z_{n}} \bigg) \exp \bigg( \frac{z}{z_{n}} \bigg)
\end{equation}
where `$m$' is the multiplicity of the zeros (so that $m = 0$ for
a simple zero). We can now apply Eq.~(\ref{eq:17}) to the zero
energy Jost function $\mathcal{X}(s^{'})$ and the Riemann's
$\xi(s^{'})$ function, assuming, $s^{'}_{n} = \frac{1}{2} + i
\gamma^{'}_{n}$ for $\mathcal{X}(s^{'})$ and $s^{'}_{n} =
\frac{1}{2} + i \gamma_{n}$ for Riemann's $\xi(s^{'})$.

This leads to the equality [Eq.~(\ref{eq:16})]:
\begin{equation}\label{eq:18}
\begin{aligned}
e^{H} e^{L\frac{s^{'}}{m}} & \prod^{\infty}_{n=1}
\bigg( 1 - \frac{s^{'}}{ms^{'}_{n}} \bigg)
\exp \bigg( \frac{s^{'}}{ms^{'}_{n}} \bigg) \\
 & = e^{J} e^{K\frac{s^{'}}{m}} \prod^{\infty}_{n=1}
\bigg( 1 - \frac{s^{'}}{ms_{n}} \bigg)
\exp \bigg( \frac{s^{'}}{ms_{n}} \bigg).
\end{aligned}
\end{equation}

First set $s^{'} = 0$ in Eq.~(\ref{eq:18}):
\begin{equation}
\Longrightarrow \qquad \qquad H = J
\end{equation}

Next set, {\mathversion{bold}$s^{'} = m s_{n}$}, $s_{n} = \frac{1}{2} + i
\gamma_{n}$
\begin{equation}\label{eq:20}
\begin{aligned}
\Longrightarrow \qquad e^{Ls_{n}} & \prod^{\infty}_{n=1}
\bigg( 1 - \frac{s_{n}}{s^{'}_{n}} \bigg)
\exp \bigg( \frac{s_{n}}{s^{'}_{n}} \bigg) \\
&= e^{Ks_{n}} \prod^{\infty}_{n=1} (1 - 1) \exp(1).
\end{aligned}
\end{equation}

Since RHS of Eq.~(\ref{eq:20}) \textbf{vanishes}, we conclude that
$s_{n}$ must equal $s^{'}_{n}$:
\begin{equation}
\therefore \gamma_{n} = \gamma^{'}_{n}
\end{equation}

This \textbf{establishes the key result} that we set out to prove:
\textbf{The Riemann zeros on the critical line do coincide precisely with
the zeros of the zero energy} {\mathversion{bold}$S$} \textbf{wave Jost function for}
{\mathversion{bold}$V(r) = \frac{\lambda}{r^{2}} \text{ } (\lambda >
0)$} \cite{Footnote}.

\renewcommand\refname{References and Footnotes:}

\begin{center}
\noindent \small{\textbf{Acknowledgement}}
\end{center}

\vspace{-12pt}
I am grateful to Irina Long for
her generous help in the preparation of this manuscript.

\end{document}